OpenStudy (anonymous):

Does anyone know how to find the circumference? The top of the ice-cream cone has 2.25 inch- diameter. What is the circumference? Use 3.14 for pi.

OpenStudy (dape):

Pi is simply the ratio of the circumference to the diamater, what that means is \[ \pi=\frac{C}{d} \] (C is circumference, d is diameter) or, likewise \[ \pi d=C \] So the circumference for your ice-cream cone is simply \[ \pi d=3.14*2.25=7.065=C \] So the circumference is 7.065 inches.

OpenStudy (anonymous):

Thank you! :)